# Attention-Guided Perturbation for Unsupervised Image Anomaly Detection

### v2

@@ 3.1.2 Reconstruction Branch
Feature Extractor. Given an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, we feed it into a pretrained network $\phi(\cdot)$ and only take a subset of $L$ layer outputs. Any off-the-shelf pre-trained ViT-like model, which relies on attention mechanisms, can readily serve as the feature extractor. In this paper, we follow previous work Reiss et al. (2022) and adopt DINO Caron et al. (2021) as our backbone, which is trained via self-distillation without labels, has demonstrated its efficacy in generating superior representations for VAD. Note that DINO keeps the resolution fixed along the layer hierarchy, and all the feature maps share the same dimension after packing them into 3D feature maps. We denote $F_{l}\in\mathbb{R}^{H_{f}\times W_{f}\times C_{f}}$ as the $l$-th output, and $\sigma(\cdot)$ as the layer normalization function. The representations from specific layers will undergo layer normalization individually, followed by summation:
@@
Another important reason to utilize DINO Caron et al. (2021) is that it is a vison-transformer-based architecture and its attention weights can be leveraged to generate the attention mask to guide the decoder training with minimal efforts. Therefore, in addition to the $L$ outputs, we propose to reuse the associated attention maps. denoted by $A_{l}\in\mathbb{R}^{H_{l}\times W_{l}}$, to provide the important prior cues for the image to guide the training of the decoder. The produced attention weights are taken out and will further be forwarded to the attention-guided perturbation branch for further integration.
Decoder. The decoder aims to decode the pretrained features of normal data in the training set. However, without proper design, it easily falls into the “identity shortcut” issue, especially under the multi-class setting. Recently UniAD You et al. (2022) attempts to alleviate this issue with a learnable-query decoder that incorporates additional learnable queries at each transformer layer. This design incurs additional costs. In our framework, we simply adopt a plain vision-transformer-based decoder thanks to an auxiliary attention-based perturbation branch. During training, the decoder is guided to decode the attention-guided perturbed representations derived from the feature extractor, rather than solely relying on the normal representation. Consequently, the “identity shotcut” issue will be greatly alleviated. Furthermore, our decoder has a more lightweight structure with only four layers, demonstrating strong empirical performance during inference.
3.1.3 Attention-Guided Perturbation.
The goal of the auxiliary branch is to perturb normal input for training under the guidance of attention masks. The details are described below.
Attention Mask Generation. To calculate the attention mask, we leverage the attention maps from the feature extractor and the momentum distillation of the decoder, which is formulated as below:
$\begin{array}[]{rl}A_{\text{final}}=&\Phi_{\text{norm}}(A_{\text{prior}})+\Phi_{\text{norm}}(A_{\text{learn}})\\ \textrm{where}&\\ &\left\{\begin{array}[]{l}\Phi_{\text{norm}}:=\textrm{max-min normalization function.}\\ A_{\text{prior}}=\Phi_{\text{aggr}}(\{A_{l}\}_{l=1}^{L})\\ A_{\text{learn}}=\Phi_{\text{aggr}}(\{A_{k}\}_{k=1}^{K})\\ \end{array}\right.\end{array}$ (2)
As discussed earlier, the attention weights from the feature extractor directly reflect the importance of each localization in the feature maps Zeiler and Fergus (2014). Therefore, based on the extracted subset of attention weights $\{A_{l}\}_{l=1}^{L}$, we compute the prior attention mask by $A_{\text{prior}}=\Phi_{\text{aggr}}(\{A_{l}\}_{l=1}^{L})$, where $\Phi_{aggr}$ denotes the aggregation operation over the attention maps. For simplicity, we only apply element-wise average pooling on the attention maps.
Second, the attention weights within the decoder are learned to aggregate important contextual cues for reconstruction. Hence, we further propose to utilize it for attention mask generation. However, these attention weights fluctuate rapidly with high variance, especially at the beginning of the training. For stabilization, we employ mean-distillation Tarvainen and Valpola (2018), which includes an exponential-moving-average (EMA) version of the decoder as the teacher.
@@
Attention-Guided Noise. For unsupervised anomaly detection, synthesized anomalies play a critical role in learning a compact and discriminative boundary to discern normal and abnormal samples. Numerous approaches have been proposed Zavrtanik et al. (2021); Li et al. (2021) for synthesizing anomalies that closely resemble real ones. Concerning the impracticality of synthesizing all types of real anomalies, we argue that focusing on critical locations and perturbing them at the feature level could be a more efficient strategy. Besides, by perturbing those critical and informative locations, the models are forced to reason the relations across different localizations more comprehensively. To this end, we simply add Gaussian noise weighted by the attention mask in the normal feature space to synthesize the hard abnormal samples. Specifically, a noise tensor $\mathcal{E}\in\mathbb{R}^{H_{f}\times\ W_{f}\times C_{f}}$ is first generated with each entry simply drawn from an i.i.d Gaussian distribution $\mathcal{N}(\mu,\sigma)$. Then we add the noise tensor to the normal features $F_{\text{clean}}$ for perturbation based on attention mask $A_{\text{final}}$ as follows:
@@
where $\gamma$ represents the basic noise factor, $p$ represents the maximum noise intensity, $m$ represents the minimum noise intensity, $T$ represents the maximun training epoch and $t$ represents the current training epoch.
@@ 3.2 Loss Function
During training, for each normal image $I_{normal}\in\mathcal{X}_{train}$, we aim to reconstruct its normal features $F_{\text{clean}}$ from for the perturbed counterparts at both the image and feature levels. Our total loss is derived as
@@
{$L_{\text{feat}}&=\frac{1}{H_{f}\times W_{f}}\text{MSE}(F_{\text{feat}},F_{\text{clean}})\\ L_{\text{img,feat}}&=\frac{1}{H_{f}\times W_{f}}\text{MSE}(F_{\text{img,feat}},F_{\text{clean}})\end{aligned}\end{cases}$ (7)
@@ 4.1 Experimental Setups
Datasets. MVTec-AD Bergmann et al. (2019) is one of the most widely used industrial anomaly detection datasets. It has a training set consisting of 3629 normal images and a test set of 467/1,258 normal/anomaly images. All the images are divided into 15 categories. In this paper, we conduct experiments with two settings. One involves training a single model for all classes (multi-class setting), while the other involves training a separate model for each class (one-class setting). The former is a more efficient yet challenging task.
@@
VisA Zou et al. (2022) is another challenging anomaly detection dataset. It has a total of 10,821 images divided into 12 different categories. The training set has 8,659 normal images, and the test set includes 962/1,200 normal/anomaly images, providing a comprehensive benchmark for evaluating detection performance.
@@
MVTec-3D Bergmann et al. (2021) is a collection of 4,147 scans obtained through the industrial 3D sensor. It covers 10 different categories, each accompanied by both RGB images and corresponding 3D point cloud data. The training set consists of 2,656 images that include only anomaly-free samples. The test set comprises 1,197 images, which encompass both normal and anomalous samples. In our experiments conducted, only the RGB images were utilized, with the 3D point cloud data not being considered.
@@
Evaluation Metrics. Referring to prior research You et al. (2022); Zavrtanik et al. (2021); Defard et al. (2021), we report the anomaly detection and localization performance with Image-level and Pixel-level Area Under the Receiver Operating Curve, denoted by I-AUC and P-AUC respectively. We also adopt Per-Region-OverlapBergmann et al. (2020) to better evaluate model’s capability, denoted by PRO.
@@
Implement Details. Each image is resized to $224\times 224$. The feature extractor is initialized by pre trained DINO(ViT-S-16) Caron et al. (2021), which is frozen during training. We employ AdamWLoshchilov and Hutter (2019) optimizer with a weight decay of $1\times 10^{-4}$. We train our framework for 500 epochs on a single GPU (NVIDIA RTX 4090) with a batch size of 32. The learning rate is initially set to $1\times 10^{-3}$ and reduced by a factor of 0.1 after 200 epochs. The ratio of added noise at the image level linearly increases from 0.6 to 1.0 from the 100th to 400th epochs. The intensity of added noise $\alpha$ at the feature level linearly increases from 0 to 1.0 from the 1st to 400th epochs. The parameter $\eta$ in EMA is set to 0.9999, and the mean teacher is updated every 10 steps. In particular, we performed a 32-fold data augmentation (rotate and flip) under a few-shot setting.
@@
Table 1: Comparison of image-level/pixel-level results on MVTec AD under multi-class setting. denotes the reproduced results using the official code. Bold and underline indicate the best and the second best, respectively.
@@
ViTAD
(arXiv 2023)
MambaAD
(NeurIPS 2024)
@@
Bottle 97.5 / 87.6 / 80.7 97.7 / 91.2 / 90.6 99.7 / 98.1 / 93.2 99.7 / 98.4 / - 100.0 / 98.8 / 94.3 100.0/98.9/96.1 100.0 / 99.1 / 95.0
Cable 57.8 / 71.3 / 40.1 87.6 / 88.1 / 85.4 95.2 / 97.3 / 85.7 94.8 / 96.8 / - 98.5 / 96.2 / 90.2 98.9/96.2/90.3 99.7 / 98.0 / 92.9
Capsule 65.3 / 50.5 / 27.3 78.3 / 89.7 / 84.5 86.9 / 98.5 /89.7 89.0 / 97.1 / - 95.4 / 98.3 / 92.0 94.5/98.5/93.9 93.8 / 98.5 / 91.9
Hazelnut 93.7 / 96.9 / 78.7 99.2 / 95.7 / 87.4 99.8 / 98.1 / 92.9 99.5 / 98.3 / - 99.8 / 99.0 / 95.2 100.0/99.0/95.8 100.0 / 98.9 / 93.6
Metal_nut 72.8 / 62.2 / 66.4 85.1 / 90.9 / 85.2 99.2 / 94.8 / 84.7 99.1 / 97.3 / - 99.7 / 96.4 / 92.4 99.9/96.9/94.0 100.0 / 95.9 / 92.6
Pill 82.2 / 94.4 / 53.9 78.3 / 89.7 / 81.9 93.7 / 95.0 / 94.7 95.7 / 95.7 / - 96.2 / 98.7 / 95.3 97.5/97.6/96.2 97.6 / 98.2 / 96.2
Screw 92.0 / 95.5 / 55.2 45.5 / 93.7 / 84.0 87.5 / 98.3 / 94.9 90.7 / 97.9 / - 91.3 / 99.0 / 93.5 95.9/99.5/97.3 91.8 / 99.0 / 95.5
Toothbrush 90.6 / 97.7 / 68.9 94.7 / 97.5 / 87.4 94.2 / 98.4 / 88.0 99.7 / 99.0 / - 98.9 / 99.1 / 90.9 98.6/99.0/92.4 100.0 / 98.9 / 91.2
Transistor 74.8 / 64.5 / 39.0 82.0 / 86.0 / 83.2 99.8 / 97.9 / 93.3 99.8 / 95.1 / - 98.8 / 93.9 / 76.8 100.0/96.6/88.3 98.7 / 93.8 / 84.3
Zipper 98.8 / 98.3 / 91.9 99.1 / 97.0 / 90.7 95.8 / 96.8 / 91.6 95.1 / 96.2 / - 97.6 / 95.9 / 87.2 99.6/98.4/95.1 99.7 / 97.6 / 90.5
Carpet 98.0 / 98.6 / 93.1 95.9 / 92.4 / 90.6 99.8 / 98.5 / 93.6 99.4 / 98.6 / - 99.5 / 99.0/ 94.7 99.9/99.3/97.5 100.0 / 99.2 / 95.6
Grid 99.3 / 98.7 / 92.1 49.8 / 46.7 / 88.6 98.2 / 96.5 / 91.1 98.5 / 96.6 / - 99.7 / 98.6 /95.8 100.0/99.3/97.3 99.9 / 98.9 / 96.3
Leather 98.7 / 97.3 / 88.5 93.9 / 96.9 / 92.7 100.0 / 98.8 / 96.5 99.8 / 98.8 / - 100.0 / 99.6 / 97.9 100.0/99.4/98.8 100.0 / 97.5 / 97.6
Tile 99.8 / 98.0 / 97.0 93.7 / 93.1 / 90.6 99.3 / 91.8 / 78.6 96.8 / 92.4 / - 100.0 / 96.6 / 87.0 98.6/93.9/81.3 100.0 / 97.4 / 90.2
Wood 99.8 / 96.0 / 94.2 95.2 / 84.8 / 76.3 98.6 / 93.2 / 85.0 99.7 / 93.3 / - 98.7 / 96.4 / 88.0 99.0/94.7/92.4 98.9 / 96.6 / 90.2
Mean 88.1 / 87.2 / 71.1 85.1 / 88.9 / 86.5 96.5 / 96.8 / 90.2 97.2 / 96.8 / 90.7 98.3 / 97.7 / 91.4 98.6/97.7/93.1 98.7 / 98.0 / 92.9
Figure 3: Qualitative illustration on MVTec-AD dataset.
4.2 Anomaly Detection on MVTec-AD
Baseline. We comprehensively evaluated representative methods in both one-class and multi-class settings. In the multi-class setting, we first compare our method with two representative one-class approaches: DRAEM Zavrtanik et al. (2021) and SimpleNet Liu et al. (2023), and then compare with methods specifically designed for the multi-class setting, including UniAD You et al. (2022), RD4AD Deng and Li (2022), DiADHe et al. (2024), ViTAD Zhang et al. (2023), MambaAD He et al. (2024b). Note that all methods are trained under the multi-class scenario. In the one-class setting, the compared models includes: RD4AD Deng and Li (2022), PatchCore Roth et al. (2022), DRAEM Zavrtanik et al. (2021), DeSTSeg Zhang et al. (2023), Simplenet Liu et al. (2023).
Evaluation on MVTec-AD under the multi-class setting. As shown in Table 1, we conducted comprehensive comparative experiments on the MVTec-AD to demonstrate the superiority of our method. It can be observed that applying one-class methods to the multi-class scenario reflects undesirable performance. Furthermore, compared with existing multi-class methods, we outperform all of them, achieving the best 98.7% and 98.0% I-AUC and P-AUC, respectively. ViTAD is a competitive approach, while we outperform it in 12 out of 15 categories in detection capacity, and exceed it in 10 categories for localization performance. In the more challenging PRO metric, our model also achieves a great performance of 92.9%, with a 1.5% improvement over ViTAD. Compared to the MambaAD approach utilizing the latest Mamba-architecture decoder, our method achieves improvements of +0.1% in I-AUC and +0.2% in P-AUC, despite a 0.2% lower performance on the PRO metric.
The qualitative comparison in Figure 3 demonstrates the superior anomaly localization capability of our method compared to DiAD and UniAD. As shown in columns 4 and 10, our approach achieves exceptional precision in handling geometrically complex objects (toothbrushes) and textured surfaces (carpets). The heatmaps which generate by the proposed model accurately align with anomaly contours at the pixel level.

### v3

@@ 3.2.2 The detail of AGPNet
Feature Extractor. Given an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, we feed it into a pretrained network $\phi(\cdot)$ and only take a subset of $L$ layer outputs. Any off-the-shelf pre-trained ViT-like model, which relies on attention mechanisms, can readily serve as the feature extractor. In this paper, we follow previous work [52] and adopt DINO [53] as our backbone, which is trained via self-distillation without labels, has demonstrated its efficacy in generating superior representations for VAD. Note that DINO keeps the resolution fixed along the layer hierarchy, and all the feature maps share the same dimension after packing them into 3D feature maps. We denote $F_{l}\in\mathbb{R}^{H_{f}\times W_{f}\times C_{f}}$ as the $l$-th output, and $\sigma(\cdot)$ as the layer normalization function. The representations from specific layers will undergo layer normalization individually, followed by summation:
@@
Another important reason to utilize DINO [53] is that it is a vison-transformer-based architecture and its attention weights can be leveraged to generate the attention mask to guide the decoder training with minimal efforts. Therefore, in addition to the $L$ outputs, we propose to reuse the associated attention maps. denoted by $A_{l}\in\mathbb{R}^{H_{l}\times W_{l}}$, to provide the important prior cues for the image to guide the training of the decoder. The produced attention weights are taken out and will be forwarded to the attention-guided perturbation branch for integration.
Decoder. The decoder aims to decode the pretrained features of normal data in the training set. In our framework, we simply adopt a plain vision-transformer-based decoder thanks to an auxiliary attention-based perturbation branch. During training, the decoder is guided to decode the attention-guided perturbed representations derived from the feature extractor. Note that our decoder is lightweight and has only four layers, yet it demonstrates strong empirical performance.
Attention Mask Generation. The goal of the auxiliary branch is to perturb normal input for training under the guidance of attention masks. The details are described below. To calculate the attention mask, we leverage the attention maps from the feature extractor and the momentum distillation of the decoder, which is formulated as below:
[…]
As discussed earlier, the attention weights from the feature extractor directly reflect the importance of each localization in the feature maps [54]. Therefore, based on the extracted subset of attention weights $\{A_{l}\}_{l=1}^{L}$, we compute the prior attention mask by $A_{\text{prior}}=\Phi_{\text{aggr}}(\{A_{l}\}_{l=1}^{L})$, where $\Phi_{aggr}$ denotes the aggregation operation over the attention maps. For simplicity, we only apply element-wise average pooling on the attention maps.
Second, the attention weights within the decoder are learned to aggregate important contextual cues for reconstruction. Hence, we further propose to utilize it for attention mask generation. However, these attention weights fluctuate rapidly with high variance, especially at the beginning of the training. For stabilization, we employ mean-distillation [29], which includes an exponential-moving-average (EMA) version of the decoder as the teacher.
@@
Attention-Guided Noise. For unsupervised anomaly detection, synthesized anomalies play a critical role in learning a compact and discriminative boundary to discern normal and abnormal samples. Numerous approaches have been proposed [3, 32] for synthesizing anomalies that closely resemble real ones. Concerning the impracticality of synthesizing all types of real anomalies, we argue that focusing on critical locations and perturbing them at the feature level could be a more efficient strategy. Besides, by perturbing those critical and informative locations, the models are forced to reason the relations across different localizations more comprehensively. To this end, we simply add Gaussian noise weighted by the attention mask in the normal feature space to synthesize the hard abnormal samples. Specifically, a noise tensor $\mathcal{E}\in\mathbb{R}^{H_{f}\times\ W_{f}\times C_{f}}$ is first generated with each entry simply drawn from an i.i.d Gaussian distribution $\mathcal{N}(\mu,\sigma)$. Then we add the noise tensor to the normal features $F_{\text{clean}}$ for perturbation based on attention mask $A_{\text{final}}$ as follows:
@@
where $\gamma$ represents the basic noise factor, $p$ denotes the maximum noise intensity, $m$ represents the minimum noise intensity, $T$ represents the maximun training epoch and $t$ represents the current training epoch.
@@ 3.3 Loss Function
During training, for each normal image $I_{normal}\in\mathcal{X}_{train}$, we aim to reconstruct its normal features $F_{\text{clean}}$ from the perturbed counterparts at both the image and feature levels. Our total loss is derived as
@@
$\begin{cases}\begin{aligned} L_{\text{feat}}&=\frac{1}{H_{f}\times W_{f}}\text{MSE}(F_{\text{feat}},F_{\text{clean}})\\ L_{\text{img,feat}}&=\frac{1}{H_{f}\times W_{f}}\text{MSE}(F_{\text{img,feat}},F_{\text{clean}})\end{aligned}\end{cases}$ (7)
@@ 4.1 Experimental Setups
Datasets. MVTec-AD [1] is one of the most widely used industrial anomaly detection datasets. It has a training set consisting of 3629 normal images and a test set of 467/1,258 normal/anomaly images. All the images are divided into 15 categories. In this paper, we conduct experiments with two settings. One involves training a single model for all classes (multi-class setting), while the other involves training a separate model for each class (one-class setting). The former is a more efficient yet challenging task.
@@
VisA [55] is another challenging anomaly detection dataset. It has a total of 10,821 images divided into 12 different categories. The training set has 8,659 normal images, and the test set includes 962/1,200 normal/anomaly images, providing a comprehensive benchmark for evaluating detection performance.
@@
MVTec-3D [56] is a collection of 4,147 scans obtained through the industrial 3D sensor. It covers 10 different categories, each accompanied by both RGB images and corresponding 3D point cloud data. The training set consists of 2,656 images that include only anomaly-free samples. The test set comprises 1,197 images, which encompass both normal and anomalous samples. In our experiments conducted, only use the RGB images, with the 3D point cloud data not being considered.
@@
Evaluation Metrics. Referring to prior research [26, 3, 2], we report the anomaly detection and localization performance with Image-level and Pixel-level Area Under the Receiver Operating Curve, denoted by I-AUC and P-AUC respectively. We also adopt Per-Region-Overlap[57] to better evaluate model’s capability, denoted by PRO.
@@
Implement Details. Each image is resized to $224\times 224$. The feature extractor is initialized by pre trained DINO (ViT-S-16) [53], which is kept frozen during training. We employ AdamW [58] optimizer with a weight decay of $1\times 10^{-4}$. We train our framework for 500 epochs on a single GPU (NVIDIA RTX 4090) with a batch size of 32. The learning rate is initially set to $1\times 10^{-3}$ and reduced by a factor of 0.1 after 200 epochs. The ratio of added noise at the image level linearly increases from 0.6 to 1.0 from the 100th to 400th epochs. The intensity of added noise $\alpha$ at the feature level linearly increases from 0 to 1.0 from the 1st to 400th epochs. The hyper-parameter $\eta$ in EMA is set to 0.9999, and the mean teacher is updated every 10 steps. Finally, for experiments under the few-shot setting, we perform a 32-fold data augmentation.
@@
Table 1: Comparison of results on MVTec AD, VisA, and MVTec-3D (left: multi-class, right: one-class). : official code results. Bold/underline: best/second best.
Multi-class Setting One-class Setting
Method MVTec AD VisA MVTec-3D Method MVTec AD
I-AUC P-AUC PRO I-AUC P-AUC PRO I-AUC P-AUC PRO I-AUC P-AUC PRO
DRÆM*[3] 88.1 87.2 71.1 79.1 91.3 59.1 65.2 93.2 55.0 DRÆM[3] 98.0 97.3
SimpleNet*[12] 85.1 88.9 86.5 87.2 96.8 81.4 RD4AD [7] 98.5 97.8 93.9
UniAD [26] 96.5 96.8 90.2 85.5 95.9 75.6 78.9 96.5 88.1 PatchCore[35] 99.1 98.1 93.5
DiAD[22] 97.2 96.8 90.7 86.8 96.0 75.2 84.6 96.4 87.8 DeSTSeg[6] 98.6 97.9
MSTAD[59] 98.6 97.2 - - - - - - - SimpleNet[12] 99.6 98.1
MoEAD[47] 97.7 97.0 - 93.1 98.7 - - - - MSTAD[59] 98.3 97.5
ViTAD[60] 98.3 97.7 91.4 90.5 98.2 85.1 78.7 98.0 91.3 GLAD[61] 99.3 98.6 -
Ours 98.7 98.0 92.9 92.3 98.4 85.5 84.9 98.0 92.9 Ours 99.2 98.3 93.1
Figure 3: Qualitative illustration on MVTec-AD dataset.
4.2 Anomaly Detection on MVTec-AD
Baseline. We comprehensively evaluated representative methods in both one-class and multi-class settings on MVTec-AD dataset. In the multi-class setting, we first compare our method with two representative one-class approaches: DRAEM[3] and SimpleNet[12], and then compare with methods specifically designed for the multi-class setting, including UniAD[26], RD4AD[7], DiAD[22], ViTAD[60], MoEAD[47] and MSTAD[59]. Note that all methods are trained under the multi-class scenario. In the one-class setting, the compared models includes: RD4AD[7], PatchCore[35], DRAEM[3], DeSTSeg[6], Simplenet[12], GLAD[61] and MSTAD[59].
Evaluation on MVTec-AD under the multi-class setting. As shown in Table 1, we conducted comprehensive comparative experiments on the MVTec-AD to demonstrate the superiority of our method. It can be observed that applying one-class methods to the multi-class scenario reflects undesirable performance. Furthermore, compared with existing multi-class methods, we outperform all of them, achieving the best 98.7% and 98.0% I-AUC and P-AUC, respectively. Our method outperforms the competitive ViTAD in both I-AUC(+0.4%) and P-AUC(+0.3%). In the more challenging PRO metric, our model also achieves a great performance of 92.9%, with a 1.5% improvement over ViTAD. For comprehensive comparisons on each category, we note our method has the following compelling advantages:
  • 1.
    Texture Categories: As shown in Table 2, existing methods already achieve near-saturation performance on texture categories. For instance, in the multi-class setting, our method attains 100.0/99.2/95.6 on Carpet and 99.9/98.9/96.3 on Grid, with only marginal improvements over ViTAD (99.5/99.0/94.7 for Carpet).
  • 2.
    Object Categories: Significant improvements are observed on object categories, particularly those with small or fine-grained anomalies (e.g., Screw: +5.6% PRO over ViTAD; Transistor: +7.5% PRO). As evidenced by Screw’s P-AUC/PRO (99.0%/95.5%), our method better preserves defect details compared to ViTAD (99.0%/93.5%).
Table 2: Anomaly detection results on MVTec AD under multi-class setting. denotes the reproduced results using the official code. Bold and underline indicate the best and the second best, respectively.
@@ 4.2 Anomaly Detection on MVTec-AD
ViTAD [60]
(CUIV 2025)
@@
Bottle 97.5 / 87.6 / 80.7 97.7 / 91.2 / 90.6 99.7 / 98.1 / 93.2 99.7 / 98.4 / - 100.0 / 98.8 / 94.3 100.0 / 99.1 / 95.0
Cable 57.8 / 71.3 / 40.1 87.6 / 88.1 / 85.4 95.2 / 97.3 85.7 94.8 / 96.8 / - 98.5 / 96.2 / 90.2 99.7 / 98.0 / 92.9
Capsule 65.3 / 50.5 / 27.3 78.3 / 89.7 / 84.5 86.9 / 98.5 89.7 89.0 / 97.1 / - 95.4 / 98.3 / 92.0 93.8 / 98.5 / 91.9
Hazelnut 93.7 / 96.9 / 78.7 99.2 / 95.7 / 87.4 99.8 / 98.1 / 92.9 99.5 / 98.3 / - 99.8 / 99.0 / 95.2 100.0 / 98.9 / 93.6
Metal_nut 72.8 / 62.2 / 66.4 85.1 / 90.9 / 85.2 99.2 / 94.8 / 84.7 99.1 / 97.3 / - 99.7 / 96.4 / 92.4 100.0 / 95.9 / 92.6
Pill 82.2 / 94.4 / 53.9 78.3 / 89.7 / 81.9 93.7 / 95.0 / 94.7 95.7 / 95.7 / - 96.2 / 98.7 / 95.3 97.6 / 98.2 / 96.2
Screw 92.0 / 95.5 / 55.2 45.5 / 93.7 / 84.0 87.5 / 98.3 / 94.9 90.7 / 97.9 / - 91.3 / 99.0 / 93.5 91.8 / 99.0 / 95.5
Toothbrush 90.6 / 97.7 / 68.9 94.7 / 97.5 / 87.4 94.2 / 98.4 / 88.0 99.7 / 99.0 / - 98.9 / 99.1 90.9 100.0 / 98.9 / 91.2
Transistor 74.8 / 64.5 / 39.0 82.0 / 86.0 / 83.2 99.8 / 97.9 / 93.3 99.8 / 95.1 / - 98.8 / 93.9 / 76.8 98.7 / 93.8 / 84.3
Zipper 98.8 / 98.3 / 91.9 99.1 / 97.0 / 90.7 95.8 / 96.8 / 91.6 95.1 / 96.2 / - 97.6 / 95.9 / 87.2 99.7 / 97.6 / 90.5
Carpet 98.0 / 98.6 / 93.1 95.9 / 92.4 / 90.6 99.8 / 98.5 / 93.6 99.4 / 98.6 / - 99.5 / 99.0 / 94.7 100.0 / 99.2 / 95.6
Grid 99.3 / 98.7 / 92.1 49.8 / 46.7 / 88.6 98.2 / 96.5 / 91.1 98.5 / 96.6 / - 99.7 / 98.6 / 95.8 99.9 / 98.9 / 96.3
Leather 98.7 / 97.3 / 88.5 93.9 / 96.9 / 92.7 100.0 / 98.8 / 96.5 99.8 / 98.8 / - 100.0 / 99.6 / 97.9 100.0 / 97.5 / 97.6
Tile 99.8 / 98.0 / 97.0 93.7 / 93.1 / 90.6 99.3 / 91.8 / 78.6 96.8 / 92.4 / - 100.0 / 96.6 / 87.0 100.0 / 97.4 / 90.2
Wood 99.8 / 96.0 / 94.2 95.2 / 84.8 / 76.3 98.6 / 93.2 / 85.0 99.7 / 93.3 / - 98.7 / 96.4 / 88.0 98.9 / 96.6 / 90.2
Mean 88.1 / 87.2 / 71.1 85.1 / 88.9 / 86.5 96.5 / 96.8 / 90.2 97.2 / 96.8 / 90.7 98.3 / 97.7 / 91.4 98.7 / 98.0 / 92.9
The qualitative comparison in Figure 3 demonstrates the superior anomaly localization capability of our method compared to DiAD and UniAD. As shown in columns 4 and 10, our approach achieves exceptional precision in handling geometrically complex objects (toothbrushes) and textured surfaces (carpets). The heatmaps which generate by our model accurately align with anomaly contours at the pixel level. For more qualitative result, Figure 4 illustrates our method’s localization performance on the MVTec-AD dataset, where the first six columns represent object categories and the last three columns show texture categories. Our model demonstrates excellent performance on both types of anomalies. For challenging micro-cracks (7th row, 7th column), the predicted heatmap (7th row, 9th column) achieves precise anomaly localization, particularly in capturing subtle structural defects. When dealing with damages that are difficult to distinguish from background patterns (10th row, 7th column), our model’s heatmap achieves pixel-level registration with ground truth masks. For more challenging object cases, the attention-guided perturbation network not only accurately localizes object deformations (6th row, 1st column) but also effectively identifies fine cracks in foreground objects (1st row, 4th column).
The comprehensive qualitative results further reveal that our model maintains robust performance on different categories. These visual results, combined with detailed case studies, fully demonstrate the competitive performance of our proposed framework across all MVTec-AD categories, complementing and reinforcing our quantitative evaluation.
Evaluation on MVTec-AD under the one-class setting. As shown in Table 1, we can observe that compared with the two synthetic data-based methods, DRAEM [3] and DeSTSeg [35], we outperform them by a large margin. Specifically, compared with DRAEM, we outperform it by 1.2% and 1.0% in I-AUC and P-AUC, respectively. And compared to DeSTSeg, we also bring gains of 0.6% I-AUC and 0.4% P-AUC. Moreover, we achieved the best localization result of 98.3% P-AUC, while the detection performance is slightly lower than SimpleNet, ranking second among previous arts. It should be noted that when SimpleNet is adapted to the multi-class setup, its performance significantly drops, as evidenced in Table 1. In contrast, our method demonstrates robust performance across both setups, highlighting its strong generalization capability and suitability for real-world applications. Similarly, compared to the existing GLAD [61], which possesses the best localization performance, the proposed method remains competitive, achieving a P-AUC score just 0.3% lower.